OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

Let the integers be x + (x+1) + (x + 2) x + (x+1) + (x + 2) = 267 3x + 3 = 267 3x = 264 x = 88 88+89+90=267 Largest integer=90 Another method 267/3 = 89 It's the average,so it has to be the middle one. Thus 89, then I took 89 - 1 = 88 and 89 + 1 = 90 and just put them in order. 88+89+90=267 Largest integer=90

OpenStudy (anonymous):

so the answer is 90

OpenStudy (anonymous):

yes @bubbasangster